# Microwave Cavity Mode Optimisation by Background Anti-Resonance Tuning

### v1

@@ Abstract
To derive the best oscillator phase noise when implementing a high-Q resonator, the resonant spectral line-shape must have high contrast and symmetry. Ideally this line-shape is second-order and Lorentzian, however, in a high mode density spectral region, low-Q background spurious modes interact and distort the resonance. For a sapphire-loaded cavity resonator operating with whispering gallery modes confined within the sapphire crystal, we show that this high contrast and symmetry can be achieved by meticulously changing the dimensions of the surrounding metallic cavity shield to tune the background low-Q structures into anti-resonance. This works because the high-Q resonances are primarily defined by the sapphire while the background modes are defined by the cavity shield. Alternatively, it was shown that a similar result can be achieved by exciting the high-Q resonator with a balanced microwave dipole probe in a Mach Zehnder interferometric configuration. The probe was constructed from two separate coaxial electric field probes symmetrically inserted into a cylindrical cavity resonator, from opposite sides with a small gap between them, so they can behave like an active wire dipole antenna. The power into the two separate probes may be matched with an external variable attenuator in one of the arms of the interferometer. Conversely, the phase between the two electric field probes may be changed with an external variable phase shifter, which changes the nature of the field components the probe couples to. The probe couples to the high-Q resonant modes as well as low-Q background modes, which can be made resonant or anti-resonant with respect to the high-Q modes by changing this external phase. When the background modes are in anti-resonance the line shape of the high-Q mode can be made symmetric and with higher contrast. This technique was applied to both whispering gallery sapphire modes, as well as hollow cavity resonators, without changing the dimensions of the cavity.
@@ I Introduction
High precision low phase noise frequency sources are extremely important in a range of modern applications, from fundamental tests of physics [1, 2, 3, 4, 5, 6, 7], high performance measurements [8], digital communications, navigation, radar technology and audio processing [9]. Amongst the best are the low noise sapphire oscillator both cryogenic, (which implements Pound frequency stabilization) [10, 11, 12, 13] and room temperature, (which implements interferometric signal processing for frequency stabilization)[14, 15, 16, 17]. Low phase noise resonator-oscillators operate optimally when configured with high-Q resonators with perfectly symmetric line-shape, usually of Lorentzian form, so the phase noise will be filtered optimally outside the bandwidth of the resonator [18]. It has been shown that any asymmetry, like a Fano type asymmetry [19], will cause phase noise of an oscillator to be enhanced [18, 7].
@@
In this work we investigate ways for an operational high-Q mode to exist in a background anti-resonance. In the first case we examine a cryogenic sapphire-loaded cavity resonator as shown in Fig. 1, which operates with Whispering Gallery (WG) modes within the sapphire. These resonators act as the frequency determination for the highly stable cryogenic sapphire oscillator, which may excite an atomic clock at the quantum projection noise limit [23]. For this case we succeed in realising the operation in a background anti-resonance by modifying the dimensions of the copper cavity that surrounds it, and show that it optimises the contrast and Q-factor of the operational mode. Following this we introduce the concept of a dipole antenna probe to excite a cavity mode, and excite the two arms of the antenna with a varying relative phase and amplitude within the probing mechanism. The configuration is similar to a Mach Zehnder interferometer [24], and we show how it can be implemented to improve microwave line shape symmetry and contrast simply through varying the external parameters.
Figure 1: The cryogenic sapphire-loaded cavity resonator. In this work the height, $h$, of the cavity is varied, to vary the background anti-resonances with $\Delta h=0$ corresponding to $h=52.10$ mm.
@@ II Quantifying Symmetry and Contrast
Due to the importance of high contrast and symmetry for optimising phase-noise performance, in this section the exact process of quantifying these values is described.
The contrast values presented in this section, and throughout this paper are calculated by fitting a parabola to the background of any narrow span spectra acquired, where resonant modes or anti-resonance features are removed from the dataset. The maximum dB point of the experimental mode then has the corresponding value of the background at the same frequency subtracted off to provide an estimated value of contrast.
@@ III Optimizing Contrast and Symmetry of a Cryogenic Sapphire Loaded Cavity
We investigate background anti-resonance tuning by changing the shield dimensions of a high-Q sapphire WG mode resonator as shown in Fig. 1. High quality-factor modes exist amongst background resonances and anti-resonances. The key feature in this case is that the background anti-resonances and resonances are tuned by varying the cavity height, $h$, dimension as shown in Fig. 2, whilst the high-Q resonances depend only on the crystal dimensions, as they are overwhelmingly located within the dielectric.
Figure 2: Effect of changing the cavity shield dimensions on a high-Q sapphire fundamental quasi-TM WGH16,0,0 mode centred at 11.202 GHz. Left, a $10$ $MHz$ span showing the properties of the background mode with contrast. Right, a $10$ kHz span showing the properties of the high-Q mode. a) A cylindrical shielding cavity height of $h\approx 52.10mm$, b) $\Delta h\approx 60\mu m$, c) $\Delta h\approx 70\mu m$, and d) a further $\Delta h\approx 30\mu m.$ The bottom curve shows unambiguously the best symmetry and contrast occurs in a background anti-resonance. Calculating the Q-factors from Fano fits to the transmission curve, we calculate a maximum loaded Q-factor of $2.5\times 10^{8}$ in anti-resonance, however the value only deviated below this by no more than 20 percent across the anti-resonance tuning range.
With each sequential change in the cavity height a frequency shift in the relatively broad background modes was observed, which are of the order of 30 kHz per $\mu m$ for the change in cavity height. In this case, it is therefore possible to shift an anti-resonance directly onto the observed mode. As the anti-resonance is tuned over the high-Q mode, the two are coupled. This results in the distortion of the latter and formation of Fano-resonances, as seen in Fig. 2. A large change in both contrast and symmetry can be achieved through this mechanism. This result was reproduced with a second sapphire resonator at a similar anti-resonance position, with the sapphire’s orientation chosen to maximise probe coupling to one of the doublets and minimise it to the other. The small finite coupling to the second under coupled mode causes some slight variations to the effect. These two resonators are being developed to create high-stability and low noise cryogenic sapphire oscillators [18, 25], so it is important to gain good symmetry and contrast in order to achieve optimum phase noise performance.
@@ IV Wire Dipole Probe
Figure 5: Top, photograph of the microwave interferometeric circuit. Bottom, a schematic illustrating the components: a) power splitter, b) variable attenuator, c) isolator, d) phase shifter, e) cavity resonator and f) dipole wire probes. The setup is similar to a Mach Zehnder interferometer, except that it measures resonance and anti-resonance instead of bright and dark ports. For network analyser experiments port 1 is labelled at the input of the interferometer, whilst probe 2 is labelled as the output probe on the side of the cavity.
@@ V Microwave dipole probe interferometer
To use such a wire dipole antenna to excite an empty cylindrical cavity resonator, it was set up symmetrically about the centre of the cavity, as shown in Fig. 4, with so-called top and bottom antennae as depicted. Control over the phase and attenuation difference between the top and bottom antennae can be achieved by use of elements similar to a Mach Zehnder interferometer [24] as shown in Fig. 5.
@@
As can be seen in Fig. 6, the introduction of a phase shift between the two dipole arms affects the magnitude and orientation of the electric field throughout the rest of the cavity. In turn, this affects both the coupling of the dipole probe to different cavity modes, dependent on the mode’s polarisation, as well as the extent of non-resonant cross-talk [31, 32] between the dipole probe and the output side probe. Anti-resonances will form in such a system due to destructive interference between resonant modes and either the input driving field or the antenna cross-talk, therefore tuning the phase shift across the dipole probe can effectively change the spectral location and extent of this destructive interference [33]. So in contrast to the first method of tuning anti-resonances by cavity height, this novel cavity excitation method allows background anti-resonances to be tuned by the non-destructive method of varying the phase difference between the input probe arms.
@@ VI Dipole Interferometric Antenna: Cavity Mode Excitation
VI.1 Tuning Low-Q Anti-Resonance
Fig. 9, shows the tuning of a broad anti-resonance by varying ($\Delta\phi$), with the effect of the anti-resonance’s in proximity to the 6.09 GHz mode shown at four unique positions.
Figure 9: A broad 5 GHz span (left) and narrow 50 MHz span (right) of the 6.09 GHz mode transmission spectra at varying phase shift and constant attenuation. The mode is displayed when the anti-resonance is positioned: a) to the right of mode, b) centrally on the mode, and c) to the left of the mode. The anti-resonance positions are indicated by a red arrow. The loaded Q, contrast and symmetry values are given for the respective anti-resonance positions.
@@ VI.1 Tuning Low-Q Anti-Resonance
VI.2 Tuning High-Q Anti-Resonance
Among the broad tuning background anti-resonances, there is also the presence of a few sharp or deeper anti-resonances, similar to that observed in the cryogenic sapphire resonator in the prior section. It was possible to tune one of these sharp anti-resonances to the same frequency of the 10.14 GHz mode. This led to an optimisation of contrast and symmetry shown in Fig. 10.
Figure 10: The optimisation of the 10 GHz mode when centred on antiresonance. The contrast (red, solid) and degree of symmetry, $\Sigma$ (blue, squares) is plotted on the same axes. Peak symmetry and contrast coincide near 3 mm phase shift. The two local maximum transmission spectra are shown with their corresponding locations. The spectra positioned in anti-resonance has $\approx$3 times better contrast.
The contrast and symmetry are both maximised when the sharp anti-resonance is positioned directly on the resonant mode centre near 3 mm phase shift in this case, noting that the peak here is possibly underestimated due to the limited data resolution around this applied phase shift of the interferometer. The results show an increase of over 18 dB in contrast at the second point of maximum symmetry. Other high symmetry peaks are also present in Fig. 10 as expected, which occurs when the background resonance is tuned to the 10.14 GHz mode, but in such situations the contrast is poor due to the interference with the background resonance.
@@ VII Dipole Interferometric Antenna: Sapphire Whispering Gallery Mode Excitation
In this section we show results of when the dipole antenna was coupled to a high-Q sapphire whispering gallery cavity as shown in Fig. 11, with similar supporting interferometric circuitry. In this higher Q system, several modes were examined, which displayed interesting behaviour when tuning the phase of the interferometer.
@@ VII.1 Tuning Whispering Gallery Mode Type 1
The first whispering gallery (WG) mode analysed, was centred at 9.905 GHz, as shown in Fig. 12. From COMSOL finite element modelling, we identify this mode as WGH11,1,0. The results also show that the optimal symmetry and contrast occur when the operational mode is tuned within a background anti-resonance. The sapphire mode’s contrast improved by several dB, along with the line shape symmetry when positioned on an anti-resonance. Along with these results a Q-factor increase of nearly $2\%$ is observed. The effect of the side-probe coupling on the dipole tuning features of this mode are also examined in Fig. 12.
@@ VIII Conclusion
In this paper, two techniques of tuning anti-resonances within a microwave cavity system were investigated, including a novel Mach-Zehnder configuration to excite microwave cavities with a generalized dipole antenna configuration. The centring of coupled background anti-resonances onto the modes of interest produced characteristic improvements in mode symmetry, contrast, and often loaded Q-factor. These features are desirable in many applications such as oscillator phase noise performance.
@@ Acknowledgments
This work was funded by the Australian Research Council Centre of Excellence for Engineered Quantum Systems, CE170100009 and the Australian Research Council Centre of Excellence for Dark Matter Particle Physics, CE200100008.
Thank you to Misty Lakelin for contributing the dipole antenna diagram in Fig. 3.

### v2

@@ Abstract
To derive the best oscillator phase noise when implementing a high-Q resonator, the resonant spectral line-shape must have high contrast and symmetry. Ideally this line-shape is second-order and Lorentzian, however, in a high mode density spectral region, low-Q background spurious modes interact and distort the resonance. For a sapphire-loaded cavity resonator operating with whispering gallery modes confined within the sapphire crystal, we show that this high contrast and symmetry can be achieved by meticulously changing the dimensions of the surrounding metallic cavity shield to tune the background low-Q structures into anti-resonance. This works because the high-Q resonances are primarily defined by the sapphire while the background modes are defined by the cavity shield. Alternatively, it was shown that a similar result can be achieved by exciting the high-Q resonator with a balanced microwave dipole probe in a Mach Zehnder interferometric configuration. The probe was constructed from two separate coaxial electric field probes symmetrically inserted into a cylindrical cavity resonator, from opposite sides with a small gap between them, so they can behave like an active wire dipole antenna. The power into the two separate probes may be matched with an external variable attenuator in one of the arms of the interferometer. Conversely, the phase between the two electric field probes may be changed with an external variable phase shifter, which changes the nature of the field components the probe couples to. The probe couples to the high-Q resonant modes as well as low-Q background modes, which can be made resonant or anti-resonant with respect to the high-Q modes by changing this external phase. When the background modes are in anti-resonance the line shape of the high-Q mode can be made symmetric and with higher contrast. This technique has been applied to both whispering gallery sapphire modes, as well as hollow-cavity resonators, without changing the dimensions of the cavity.
@@ I Introduction
High precision low phase noise frequency sources are extremely important in a range of modern applications, from fundamental tests of physics [1, 2, 3, 4, 5, 6, 7], low-noise measurements [8], through digital communications, navigation, radar technology to audio processing [9]. Amongst the best is the low noise sapphire oscillator both cryogenic, (which implements Pound frequency stabilization) [10, 11, 12, 13] and room temperature, (which implements interferometric signal processing for frequency stabilization)[14, 15, 16, 17]. Low phase noise resonator-oscillators operate optimally when configured with high-Q resonators with perfectly symmetric line-shape, usually of Lorentzian form, so the phase noise will be filtered optimally outside the bandwidth of the resonator [18]. It has been shown that any asymmetry, like a Fano type asymmetry [19], will cause phase noise of an oscillator to be enhanced [18, 7].
@@
In this work we investigate ways for an operational high-Q mode to exist in a background anti-resonance. In the first case we examine a high-Q Whispering Gallery (WG) mode sapphire-loaded cavity resonator cooled to near 4 K, as shown in Fig. 1. This type of resonator acts as the frequency determination for the highly stable cooled sapphire oscillator, which may excite an atomic clock at the quantum projection noise limit [23]. For this case we succeeded in realising the operation in a background anti-resonance by modifying the dimensions of the copper cavity that surrounds it, and show that it optimises the contrast and Q-factor of the operational mode. Following this we introduce the concept of a dipole antenna probe with two input arms to excite the cavity resonator. The probing mechanism configuration is similar to a Mach Zehnder interferometer [24], as the two arms may have the relative phase and amplitude varied. In this case we show how it can be implemented to improve microwave line shape symmetry and contrast simply through varying the external parameters. In this case we only undertook measurements at room temperature as a proof of concept experiment.
Figure 1: The cryogenic sapphire-loaded cavity resonator. In this work the height, $h$, of the cavity was varied, which tuned the background anti-resonances without tuning the sapphire mode. In this work, $\Delta h=0$ corresponded to a value of $h=52.10$ mm.
Figure 2: Effect of changing the cavity shield dimensions on a high-Q sapphire fundamental quasi-TM WGH16,0,0 mode centred at 11.202 GHz. Left, a $10$ $MHz$ span showing the properties of the background mode with contrast. Right, a $10$ kHz span showing the properties of the high-Q mode. a) A cylindrical shielding cavity height of $h\approx 52.10mm$, b) $\Delta h\approx 60\mu m$, c) $\Delta h\approx 70\mu m$, and d) a further $\Delta h\approx 30\mu m.$ The bottom curve shows unambiguously the best symmetry and contrast occurs in a background anti-resonance. Calculating the Q-factors from Fano fits to the transmission curve leads to a maximum loaded Q-factor of $2.5\times 10^{8}$ in anti-resonance, however the value only deviated below this by no more than 20 percent across the anti-resonance tuning range.
@@ II Quantifying Symmetry and Contrast
Due to the importance of high contrast and symmetry for optimising phase-noise performance, in this section the exact process of quantifying these values is described. The contrast values presented in this section, and throughout this paper are calculated by fitting a parabola to the background of any narrow span spectra acquired, where resonant modes or anti-resonance features are removed from the dataset. The maximum dB point of the experimental mode then has the corresponding value of the background at the same frequency subtracted off to provide an estimated value of contrast.
@@ III Optimizing Contrast and Symmetry of a Cryogenic Sapphire Loaded Cavity
We investigate background anti-resonance tuning by changing the shield dimensions of a high-Q sapphire WG mode resonator as shown in Fig. 1. High quality-factor modes exist amongst background resonances and anti-resonances. The key feature in this case is that the background anti-resonances and resonances are tuned by varying the cavity height, $h$ (dimension as shown in Fig. 2), whilst the high-Q resonances depend only on the crystal dimensions, as they are overwhelmingly located within the dielectric.
With each sequential change in the cavity height a frequency shift in the relatively broad background modes was observed, on the order of 30 kHz per $\mu$m for the change in cavity height. In this case, it was therefore possible to shift an anti-resonance directly onto the observed mode. As the background anti-resonance was tuned over the high-Q mode, distortion of the high-Q mode was observed, with a large change in both contrast and symmetry. This affect may be explained by the onset of a Fano-resonance, as seen in Fig. 2, modelled through a mutual resistive term between the two modes as described in [20]. This result was reproduced with a second sapphire resonator at a similar anti-resonance position, with the sapphire’s orientation chosen to maximise probe coupling to one of the doublets and minimise it to the other. The small finite coupling to the second under coupled mode causes some slight variations to the effect. These two resonators are being developed to create high-stability and low noise cryogenic sapphire oscillators [18, 25], so it is important to gain good symmetry and contrast in order to achieve optimum phase noise performance.
@@ IV Wire Dipole Probe
Figure 5: Top, photograph of the microwave interferometeric circuit. Bottom, a schematic illustrating the components: a) power splitter, b) variable attenuator, c) isolator, d) phase shifter, e) cavity resonator and f) dipole wire probes. The setup is similar to a Mach Zehnder interferometer, except that it measures resonance and anti-resonance instead of bright and dark ports. For network analyser experiments port 1 is labelled as the input of the interferometer, whilst probe 2 is labelled as the output probe on the side of the cavity.
@@ V Microwave dipole probe interferometer
To use such a wire dipole antenna to excite an empty cylindrical cavity resonator, it was set up symmetrically about the centre of the cavity, as shown in Fig. 4, with so-called top and bottom as depicted. Control over the phase and attenuation difference between the top and bottom antennas can be achieved by use of elements similar to a Mach Zehnder interferometer [24] as shown in Fig. 5.
@@
As illustrated in Fig. 6, the introduction of a phase shift between the two dipole arms affects the orientation and magnitude of the electric field that the probe couples to. This affects the coupling of the dipole probe to various cavity modes by simply changing the interferometer parameters, and is dependent on the mode’s polarisation, as well as the extent of the coupling to cross-talk [31, 32] between the dipole probe and the output side probe. Anti-resonances will form due to destructive interference between resonant modes and either the input driving field or the cross-talk, therefore tuning the phase shift across the dipole probe can effectively change the spectral location and extent of destructive and constructive interference [33]. Thus, anti-resonance and resonance frequencies may be tuned via the phase difference imposed by the interferometer. So in contrast to the first method of tuning anti-resonances by cavity height, this novel cavity excitation method allows background anti-resonances to be tuned by the non-destructive method of varying the phase difference between the input probe arms.
@@ VI Dipole Interferometric Antenna: Cavity Mode Excitation
VI.1 Tuning a Low-Q Anti-Resonance
Fig. 9, shows the tuning of a broad anti-resonance by varying ($\Delta\phi$), with the effect of the anti-resonance’s in proximity to the $TM_{1,1,0}$ 6.09 GHz mode shown at four unique positions.
Figure 9: A broad 5 GHz span (left) and narrow 50 MHz span (right) of the $TM_{1,1,0}$ 6.09 GHz mode transmission spectra at varying phase shift and constant attenuation. The mode is displayed when the anti-resonance is positioned: a) to the right of mode, b) centrally on the mode, and c) to the left of the mode. The anti-resonance positions are indicated by a red arrow. The loaded Q, contrast and symmetry values are given for the respective anti-resonance positions.
@@ VI.1 Tuning a Low-Q Anti-Resonance
VI.2 Tuning a Higher-Q Anti-Resonance
Among the broad tuning background anti-resonances, there is also the presence of a few sharp or deeper anti-resonances, similar to that observed in the cryogenic sapphire resonator in the prior section. It was possible to tune one of these sharp anti-resonances to the same frequency of the $TM_{2,1,1}$ 10.14 GHz mode. This led to an optimisation of contrast and symmetry shown in Fig. 10.
Figure 10: The optimisation of the $TM_{2,1,1}$
10.14 GHz mode when centred on antiresonance. The contrast (red, solid) and degree of symmetry, $\Sigma$ (blue, squares) is plotted using a double y-axis. Peak symmetry and contrast coincided near 3 mm phase shift. The two local maximum transmission spectra are shown with their corresponding locations. The spectra positioned in anti-resonance has $\approx$3 times better contrast. The Q-factor of the mode is independent on phase shift and equal to 4,600.
The contrast and symmetry are both maximised when the sharp anti-resonance is positioned directly on the resonant mode centre near 3 mm phase shift in this case, noting that the peak here is possibly underestimated due to the limited data resolution around this applied phase shift of the interferometer. The results show an increase of over 18 dB in contrast at the second point of maximum symmetry. Other high symmetry peaks are also present in Fig. 10 as expected, which occurs when the background resonance is tuned to the $TM_{2,1,1}$ 10.14 GHz mode, but in such situations the contrast is poor due to the interference with the background resonance.
@@ VII Dipole Interferometric Antenna: Sapphire Whispering Gallery Mode Excitation
In this section we show room temperature results of when the dipole antenna was coupled to a high-Q sapphire whispering gallery cavity as shown in Fig. 11, with similar supporting interferometric circuitry. These modes are more than an order of magnitude higher-Q than the empty copper cavity modes. In this higher Q system, several modes were examined, which displayed interesting behaviour when tuning the phase of the interferometer.
@@ VII.1 Tuning Whispering Gallery Mode Type 1
The first whispering gallery (WG) mode analysed, was centred at 9.905 GHz, as shown in Fig. 12. From COMSOL finite element modelling, we identify this mode as WGH11,1,0. The results also show that the optimal symmetry and contrast occured when the operational mode was tuned within a background anti-resonance. The sapphire mode’s contrast improved by several dB, along with the line shape symmetry when positioned on an anti-resonance. Along with these results a Q-factor increase of about $2\%$ was observed. The effect of the side-probe coupling on the dipole tuning features of this mode were also examined in Fig. 12.
@@ VIII Conclusion
In this paper, two techniques of tuning anti-resonances within a microwave cavity system were investigated, including a novel Mach-Zehnder configuration to excite microwave cavities with a generalized dipole antenna configuration. The centring of coupled background anti-resonances onto the modes of interest produced large improvements in mode symmetry and contrast. These features are desirable in many applications such as oscillator phase noise performance.
@@ Acknowledgments
This work was funded by the Australian Research Council Centre of Excellence for Engineered Quantum Systems, CE170100009 and the Australian Research Council Centre of Excellence for Dark Matter Particle Physics, CE200100008. Thank you to Misty Lakelin for contributing the dipole antenna diagram in Fig. 3.